# Guardians of the Quantum GAN

### v2

@@ V-C Watermark extraction
Extraction for Single Hardware Training: We consider that the user has trained the qGAN on a single hardware. To train the classifier, we generate a dataset from different IBM backends. We train the qGAN model on all 10 IBM backends and generate the images while inferencing on the backends. We generate 100 images every time the trained model is inference on a different IBM backend. This way we obtain a rich dataset of images having 10 labels representing the hardware where the qGAN was trained. From Fig 7 (1, 3) we can observe that the classifier trains with near-perfect accuracy. Owing to the richness of the dataset produced due to the influence of noise of the IBM backends, the task of earning patterns in the data is not a very complex task. To make the extraction task very efficient, we set a threshold value, $\mathcal{M}$. To define $\mathcal{M}$, we test the classifier with a test dataset comprising images generated by the trained qGAN on different IBM backends and set $\mathcal{M}$ as the mean of the probability of the detected training hardware by the classifier for each image. To ensure that the test and training datasets differ, we re-train the qGAN and re-generate the images using random seed values. For the situation where the user trains on a single hardware, we obtain $\mathcal{M}=0.9999$ for our qGAN model and dataset. We demonstrate a subset of the results in Fig 8, where we test the trained classifier using a test dataset consisting of 5000 images where the images generated from qGANs trained on IBM Washington and IBM Cambridge have previously not been seen by the classifier. We can observe that the images generated from known backends are classified correctly with a probability of $\sim 1$. Although the images generated from unknown sources are classified as known hardware, the probability of predicting the correct label of hardware is less than the value of $\mathcal{M}(=0.9999)$, which proves our case of handling cases where the suspected qGAN models belong to the user (known hardware) or not (unknown hardware).
@@
Extraction for Multiple Hardware Training: To demonstrate this procedure we pick 4 IBM backends from our suite having 5, 7, 27, and 127 qubits and train the qGAN on all possible combinations taking two at a time providing a set of 12 labels of known hardware. We also include the dataset of images generated from the single hardware whose combinations we provide, making a final dataset of 16,000 images generated from the qGAN model trained on 12 combinations of 4 IBM backends and the 4 single backends, to train the classifier. We observe from Fig. 7 (2,4) that the classifier trains with a reduced accuracy of around 90%. Owing to the multiple noise models influencing the generated images, the quality of the dataset used to train the classifier was reduced slightly, therefore decreasing the accuracy of the prediction. We find that while testing the trained classifier with a dataset of images generated from known combinations of hardware, the mean probability of detecting the correct hardware combination is almost equal to 1($\sim 0.9$) which becomes our threshold value, $\mathcal{M}$. We describe a subset of these results in Fig. 8 where we test the classifier on a dataset of 5000 images generated similarly as the first use case to ensure that the classifier has never seen the test data. We observe that the images generated from the qGAN trained on known hardware combinations get predicted with a probability of $\sim 1$. Although the images generated from unknown sources are classified as known hardware combinations, the probability of predicting the correct label of hardware is less than the value of $\mathcal{M}(=0.9)$, which proves our case of handling cases where the suspected qGAN models belong to the user (known hardware) or not (unknown hardware).
Figure 8: Plots describing the high predicted probability for known hardware during the proof of ownership using the classifier.
@@
We also counter a corner case where the user trains the qGAN model on a set of quantum hardware in a particular sequence, and the suspected qGAN model has been trained on the same set of quantum hardware but in a different sequence. We validate our results in this case by training the classifier on a reduced dataset from the sub-section of model migration. We train the qGAN model on unique combinations from the 4 chosen IBM backends, taking 2 at a time. We can observe from Fig. 7 that the classifier in this situation trains with almost 100% accuracy. We proceed to calculate the threshold value ($\mathcal{M}$) to be $0.999$. To test the case, we prepare a dataset comprising 4000 images with images generated by the qGAN trained on two known combinations in two different orders. Fig. 8 shows that the known hardware combinations are predicted with a near-perfect probability and the unknown sequence of hardware gets classified with a probability much less than the threshold value, $\mathcal{M}(=0.999)$.

### v3

@@ V-C Watermark extraction
Extraction for Single Hardware Training: We consider that the user has trained the qGAN on a single hardware. To train the classifier, we generate a dataset from different IBM backends. We train the qGAN model on all 10 IBM backends and generate the images while inferencing on the backends. We generate 100 images every time the trained model is inference on a different IBM backend. This way we obtain a rich dataset of images having 10 labels representing the hardware where the qGAN was trained. From Fig 7 (1, 3) we can observe that the classifier trains with near-perfect accuracy. Owing to the richness of the dataset produced due to the influence of noise of the IBM backends, the task of earning patterns in the data is not a very complex task. To make the extraction task very efficient, we set a threshold value, $\mathcal{M}$. To define $\mathcal{M}$, we test the classifier with a test dataset comprising images generated by the trained qGAN on different IBM backends and set $\mathcal{M}$ as the mean of the probability of the detected training hardware by the classifier for each image. To ensure that the test and training datasets differ, we re-train the qGAN and re-generate the images using random seed values. For the situation where the user trains on a single hardware, we obtain $\mathcal{M}=0.9999$ for our qGAN model and dataset. We demonstrate a subset of the results in Fig 8.1, where we test the trained classifier using a test dataset consisting of 5 labels, each having 1000 images. We construct the test dataset using images generated from qGANs trained on two quantum hardware in our suite (IBM Washington and IBM Cambridge), and three not included in our suite (Unknown1-IBM Casablanca, Unknown2-IBM Almaden, Unknown3-IBM Essex). We can observe that the images generated from known backends are classified correctly with a probability of $\sim 1$. Although the images generated from unknown sources are classified as known hardware, the probability of predicting the correct label of hardware is less than the value of $\mathcal{M}(=0.9999)$, which proves our case of handling cases where the suspected qGAN models belong to the user (known hardware) or not (unknown hardware).
@@
Extraction for Multiple Hardware Training: To demonstrate this procedure we pick 4 IBM backends from our suite having 5 (IBM Athens), 7 (IBM Jakarta), 27 (IBM Kolkata), and 127 (IBM Washington) qubits and train the qGAN on all possible combinations taking two at a time providing a set of 12 labels of known hardware. We also include the dataset of images generated from the single hardware whose combinations we provide, making a final dataset of 16,000 images generated from the qGAN model trained on 12 combinations of 4 IBM backends and the 4 single backends, to train the classifier. We observe from Fig. 7(2,4) that the classifier trains with a reduced accuracy of around 90%. Owing to the multiple noise models influencing the generated images, the quality of the dataset used to train the classifier was reduced slightly, therefore decreasing the accuracy of the prediction. We find that while testing the trained classifier with a dataset of images generated from known combinations of hardware, the mean probability of detecting the correct hardware combination is almost equal to 1($\sim 0.9$) which becomes our threshold value, $\mathcal{M}$. We describe a subset of these results in Fig. 8.2 where we test the classifier on a dataset having 5 labels, each with 1000 images generated similarly as the first use case to ensure that the classifier has never seen the test data. The test dataset has images generated by qGANs trained on known quantum hardware (a sequence of IBM Athens and IBM Kolkata, and a sequence of IBM Jakarta and IBM Athens), and three unknown quantum hardware (Unknown1-IBM Casablanca, Unknown2-IBM Almaden, Unknown3-IBM Essex). We observe that the images generated from the qGAN trained on known hardware combinations get predicted with a probability of $\sim 1$. Although the images generated from unknown sources are classified as known hardware combinations, the probability of predicting the correct label of hardware is less than the value of $\mathcal{M}(=0.9)$, which proves our case of handling cases where the suspected qGAN models belong to the user (known hardware) or not (unknown hardware).
Figure 8: In these plots we describe the high predicted probability for known hardware while validating the proof of ownership using the classifier. The classifier is trained to classify images generated by the qGAN trained on the hardware marked on the x-axis. Figure (1) demonstrates the validation during the training of the qGAN on a single quantum hardware, (2) demonstrates the validation during the training of the qGAN on a sequence of quantum hardware; (3) demonstrates the prominence of the sequence of hardware on which the qGAN is trained.
@@
We also counter a corner case where the user trains the qGAN model on a set of quantum hardware in a particular sequence, and the suspected qGAN model has been trained on the same set of quantum hardware but in a different sequence. We validate our results in this case by training the classifier on a reduced dataset from the sub-section of model migration. We train the qGAN model on unique combinations from the 4 chosen IBM backends, taking 2 at a time. We can observe from Fig. 7 that the classifier in this situation trains with almost 100% accuracy. We proceed to calculate the threshold value ($\mathcal{M}$) to be $0.999$. To test the case, we prepare a dataset comprising 4 labels, each with 1000 images generated by the qGAN trained on two known combinations (a sequence of IBM Athens and IBM Kolkata, and a sequence of IBM Jakarta and IBM Washington) in two unknown combinations (a sequence of IBM Kolkata and IBM Athens, and a sequence of IBM Washington and IBM Jakarta). Fig. 8.3 shows that the known hardware combinations are predicted with a near-perfect probability and the unknown sequence of hardware gets classified with a probability much less than the threshold value, $\mathcal{M}(=0.999)$. This proves that the noise watermark embedded in the images generated by the qGAN depends on the exact sequence of the hardware on which the qGAN is trained.